# GreedyML: A Parallel Algorithm for Maximizing Submodular Functions

### v1

@@ 2.3 The Greedi  and RandGreedi Algorithms
We now discuss the Greedi and RandGreedi, which are the state-of-the-art distributed algorithms for constrained submodular maximization. The Greedi algorithm [24] partitions the data arbitrarily on available machines, and on each partition, it runs the Greedy algorithm in parallel to compute a local solution. These solutions are then sent to a single global machine, where they are accumulated. The Greedy algorithm is then again executed on the accumulated data to get a global solution. The final solution is the best solution among all the local and global solutions. For a cardinality constraint, where $k$ is the solution size, the Greedi algorithm has a worst-case approximation guarantee of $1/\Theta(\min(\sqrt{k},m))$, where $m$ is the number of machines.
Although the Greedi algorithm performs well in practice [24], its approximation ratio is not a constant but depends on $k$. To improve the approximation guarantee of Greedi algorithm, Barbosa et al. proposed the RandGreedi algorithm [2]. By partitioning the data uniformly at random on machines, RandGreedi achieves an expected approximation guarantee of $\frac{1}{2}(1-1/\mathrm{e})$ for cardinality and $1/4$ for matroid constraints. In general, it has an approximation ratio of $\alpha/2$ where $\alpha$ is the approximation ratio of the Greedy algorithm used at the local and global machines. We present the pseudocode of RandGreedi framework in Algorithm 2. Note that for a cardinality constraint, both Greedi and RandGreedi perform $O(nk(k+m))$ calls to the objective function and has $O(mk)$ elements communicated to the single central machine where $n$ is the number of elements in the ground set, $m$ is the number of machines, and $k$ is solution size.
Both Greedi and RandGreedi require a single global accumulation from the solutions generated in local machines. This single accumulation step can quickly become dominating since the runtime, memory, and complexity of this global aggregation grows linearly with the number of machines. We propose to alleviate this by introducing a hierarchical aggregation strategy that maintains an accumulation tree. Our GreedyML framework generalizes the RandGreedi from a single accumulation to a multi-level accumulation. The number of partial solutions to be aggregated depends on the branching factor of the tree, which can be a constant. Thus, the number of accumulation levels grows logarithmically with the number of machines, and the total aggregation is not likely to become a memory, runtime, and communication bottleneck with the increase in the number of machines.
2.4 Other Related Work
Kumar et al. [17] have developed the sample and prune algorithm which achieves an expected approximation ratio of $1/(2+\varepsilon)$ for $k$-cardinality constraints, using $O(1/\delta)$ rounds, when the memory per machine is $O(kn^{\delta}\>\log{n})$, where $\delta>0$ is a parameter, and $n$ is the number of elements in the ground set. Barbosa et al. [2] have compared their RandGreedi algorithm with this one and show that the former performs better than the latter for the practical quality of the computed approximate solution. They observed this even though the sample and prune algorithm has a better worst-case approximation ratio in expectation.
More recent work on distributed submodular maximization uses the multi-linear extension to map the submodular optimization problem into a continuous domain. This line of work [4, 25, 26] typically performs a gradient ascent on each local machine and builds a consensus solution in each round, which improves the approximation factor to $(1-1/\mathrm{e})$. However, we believe these latter papers represent primarily a theoretical contribution rather than one that leads to practical algorithms. The reason is the high (exponential) cost of computing a single gradient by sampling many points; even randomized approximations of gradient computations are expensive. Most of these algorithms are not implemented and the ones with implementation solve problems with only a hundred elements in the data set [26].
3 Description of Our Algorithm
We describe and analyze our algorithm that generalizes the RandGreedi algorithm from a single accumulation step to multiple accumulation steps. Each accumulation step corresponds to a level in an accumulation tree which we describe next. We assume that there are $m$ machines identified by the set of ids: $\{0,1,\ldots,m-1\}$.
Accumulation Tree. An accumulation tree ($T$) is defined by the number of machines ($m$), and branching factor ($b$). It has the same structure as a complete $b$-ary tree with $m$ leaves which means all the leaves are at the same depth. The tree nodes correspond to processors along with the corresponding subset of data accessible to them. The edges of the tree determine the accumulation pattern of the intermediate solutions. The final solution is generated on the root node of $T$. Thus, the branching factor $b$ of the tree indicates the maximum number of nodes that send data to its parent. For each internal node of the tree, we attempt to have exactly $b$ children. Note that since we plan to construct a complete $b$-ary tree, in the case where $m$ is not multiple of $b$, in each level of the tree, there could be at most one node whose arity is less than $b$. The number of accumulation levels, $L$ (i.e., the height of the tree minus 1) is $\lceil\log_{b}{m}\rceil$.
To uniquely identify a node in the tree, we will assign an identifier $(\ell,id)$ to each node of $T$, where $\ell$ represents the accumulation level of the node and $id$ represents the machine id corresponding to the node. The $id$ for each leaf node is the id of the machine that the leaf node corresponds to. All the leaf nodes are at level 0. Each internal node receives the lowest $id$ of its children, i.e., any node $(l,i)$ has node $(l+1,\lfloor i/b^{l+1}\rfloor*b^{l+1})$ as the parent. Therefore the root node will always have level $L$ with $id$ value 00. Also, we characterize an accumulation tree $T$ by the triple $T(m,L,b)$, where $m$ is the number of leaves (machines), $L$ is the number of levels, and $b$ is the branching factor.
{forest}
[…]
{forest}
{forest}
{forest}
{forest}
Figure 2: Accumulation tree with 8 machines and branching factors 2 (top-left), 3 (top-right), 4 (bottom-left), and 8 (bottom-right). The level inside a node represents the identification of the node.
Figure 1 shows an example of a generic accumulation tree with $b^{2}$ leaves and branching factor $b$. The number of accumulation levels is the level of the root. Here we have $L=\lceil\log_{b}b^{2}\rceil=2$. Figure 2 shows accumulation trees with 8 leaves and with branching factors 2, 3, 4, and 8. The trees with branching factors 2 and 8 have the same branching factor for every internal node as these trees have $b^{L}$ nodes. But the tree with branching factor 3 has the last node in level 1 with only 2 children. Similarly, the tree with branching factor 4 has the last node in level 2 i.e. the root with 2 children. Observe that the $id$ parameter remains the same in multiple nodes that are involved in computations at multiple levels. For this paper, we show analysis by keeping the branching factor constant across all levels.
Data Accessibility. We use $P_{id}$ to denote the elements assigned to machine $id$. To indicate the data accessible to a particular node in the tree, we describe a set for the input data set as $V_{\ell,id}$. It corresponds to all the data used to compute the solution at the node $(\ell,id)$ and consists of all the elements assigned to its descendants:
$V_{\ell,id}=\bigcup\limits_{i=0}^{\min(b^{\ell}-1,m-id)}P_{id+i}.$
$\textsc{GreedyML}(\ell,id)=$
$\begin{cases}\textsc{Greedy}(P_{id})&\ell=0\\ \arg\max\begin{cases}\textsc{Greedy}\left(\bigcup\limits_{i\in\{0,1,\dotsc,b-1\}}\textsc{GreedyML}\left(\ell-1,id+i\cdot b^{\ell-1}\right)\right)\\ \textsc{GreedyML}(\ell-1,id)\end{cases}&id\text{ mod }b^{\ell}=0\\ \text{undefined}&\text{otherwise}\end{cases}$
Recurrence Relation. Figure 3 shows a recurrence relation defined for every node in the accumulation tree and will be the basis for our multilevel distributed algorithm. At level 00 (leaves), the recurrence function returns the Greedy solution of the random subset of data $P_{id}$ assigned to it. At other levels (internal nodes), it returns the better among the Greedy solution computed from the union of the received solution sets of its children and its solution from its previous level. It is undefined for $(\ell,id)$ tuples that do not correspond to nodes in the tree (at higher levels). We call our algorithm associated with the recurrence relation as the GreedyML algorithm.
We can compare it with the RandGreedi  algorithm by looking at the recurrence relation at level one. Our recurrence relation takes the $\arg\max$ for the accumulated solution and one solution from the previous level. However, the RandGreedi algorithm takes the $\arg\max$ of the accumulated solution and the best solution from the children. Our choice reduces the computation at the internal node. We show that this modification produces the same approximation ratio as the RandGreedi algorithm.
@@ 3 Description of Our Algorithm
Pseudocode. Algorithm 3 describes our multilevel distributed algorithm in two functions. The first function GreedyML is a wrapper function that sets up the environment to run the distributed algorithm. The second function $\textsc{GreedyML}^{\prime}$ is the iterative implementation of the recurrence relation that runs on each machine. The wrapper function partitions the data into $m$ subsets and assigns them to the machines (Line 2). Then each machine runs the $\textsc{GreedyML}^{\prime}$ function on the subset assigned to it (Line 5, Line 7). The wrapper function uses and returns the solution from machine 00 (Line 8) as it is the root of the accumulation tree.
@@
The $\textsc{GreedyML}^{\prime}$ procedure is an iterative implementation of the recurrence relation 3 that runs on every machine. Each machine checks whether it needs to be active at a particular level (Line 5) and decides whether it needs to receive from (Line 11) or send to other machines (Line 6). The function returns the solution from the last level of the machine.
@@
3:    for $i=1\ldots m-1\text{ in parallel}$ donormal-$\triangleright$ Run GreedyML’ on all machines except 00
@@
5:         Run $\textsc{GreedyML}^{\prime}(V_{i},\ell,b,i)$ to obtain $S_{i}$ on machine $i$
@@
7:    Run $\textsc{GreedyML}^{\prime}(V_{0},\lceil\log_{b}m\rceil,b,0)$ to obtain $S_{0}$ on machine 0
@@
1:procedure $\textsc{GreedyML}^{\prime}$($P$: Partial Data-set, $\ell$: levels; $b$: branching factor, $id$: machine ID)
@@
4 Analysis of Our Algorithm
In this section, we will derive the expected approximation ratio of the GreedyML algorithm. We will then describe the three submodular functions we experiment with and derive their computation and communication complexities.
4.1 Expected Approximation Ratio
This subsection proves the expected approximation ratio of our GreedyML algorithm in Theorem 4.7. To do so, we need three Lemmas. The first Lemma characterizes elements that do not change the solution computed by the GreedyML algorithm. We need some preliminary notation. When we wish to indicate the data set that a node in the tree and its descendants work with, we add an argument to $\textsc{GreedyML}(\ell,id)$, and write $\textsc{GreedyML}(V_{\ell,id},\ell,id)$. When we perform a union operation on this data set with some set $B$, and execute the GreedyML algorithm on the union, i.e., $\textsc{GreedyML}(V_{\ell,id}\cup B,\ell,id)$, then elements in $B$ are assigned randomly to the leaves of the subtree rooted at node $(\ell,id)$ and the algorithm is run with the updated data sets. Lemma 4.1 compares executions of the algorithm when this union operation is performed for a special set $B$. It states that if adding an individual element of a set, $B$, to the input of the GreedyML does not change the solution set then adding the whole set, $B$ to the input will also have no effect on the solution. Here we consider executions that use the same random tape, number of machines, and branching factor. We use the same random tape to couple the executions. Therefore, the result of the Lemma is not over the expectation of the random tape.
Lemma 4.1.
Let $T(m,L,b)$ be an accumulation tree. Consider a universal set $W$, and a random tape $r_{W}$ that maps elements of $W$ to the leaves of $T$. Let $V\subseteq W$ denote the set of elements accessible to a node $(\ell,id)$, and consider adding elements of $B\subseteq W$ to this node. If we have GreedyML $(V\cup\{e\},\ell,id)=\textsc{GreedyML}(V,\ell,id)$, for each element $e\in B$, then $\textsc{GreedyML}(V\cup B,\ell,id)=\textsc{GreedyML}(V,\ell,id)$.
Note: Function calls in this analysis use the same random tape for assigning elements; hence elements are assigned uniformly at random to the machines, but they use the same random assignment in all runs involving $V$; $V\cup\{e\},\forall e\in B$; and $V\cup B$.
Proof 4.2.
If possible, let $\textsc{GreedyML}(V\cup B,\ell,id)\neq\textsc{GreedyML}(V,\ell,id)$. Let $e$ be the first element of $B$ to be selected by the Greedy algorithm at the final level. Let $i$ be the level in which $e$ was in the input in some machine but not selected in a solution for the next level in $\textsc{GreedyML}(V\cup\{e\},\ell,id)$. Since $e$ is the first element of $B$ that was selected by the Greedy algorithm, the elements chosen before it at level $i$ in $\textsc{GreedyML}(V\cup B,\ell,id)$ are the same ones chosen before it at level $i$ in $\textsc{GreedyML}(V\cup\{e\},\ell,id)$. Since it was not selected in $\textsc{GreedyML}(V\cup\{e\},\ell,id)$ it will not be selected in $\textsc{GreedyML}(V\cup B,\ell,id)$. This is a contradiction since $e$ must be selected at every level to be present in the final solution.
Now we turn to the two Lemmas that provide bounds on the quality of the computed solutions in terms of the optimal solution at an internal node in the accumulation tree.
Lemma 4.3 provides a lower bound on the expected function value of the solution of the GreedyML algorithm from a child of the internal node. Lemma 4.5 provides a lower bound on the expected function value of the solution set from the Greedy algorithm executed at each internal node on the union of the partial solutions from its children. These Lemmas depend on the probability distribution defined below.
Let $p_{\ell,id}\colon V_{\ell,id}\to[0,1]$ be a probability distribution over the elements in $V_{\ell,id}$, which we shall define below. Here $A\sim V_{\ell,id}(1/b)$ denotes a random subset of $V_{\ell,id}$ such that each element is independently present in $A$ with probability $1/b$. This probability corresponds to the distribution from the random tape because each element is present with the same likelihood from any child of the node. Let $OPT_{\ell,id}$ be an optimal solution of the constrained submodular maximization problem when the input data is $V_{\ell,id}$.
The probability $p_{\ell,id}$ is defined as follows:
$p_{\ell,id}(e)=\begin{cases}\Pr\limits_{A\sim V_{\ell,id}(1/b)}\left[e\in\textsc{GreedyML}(A\cup\{e\},\ell-1,id)\right],&\text{if}\ e\in OPT_{\ell,id};\\ 0,&\text{otherwise.}\end{cases}$
For any internal node $(\ell,id)$, the distribution $p_{\ell,id}$ defines the probability that each element of $OPT_{\ell,id}$
is in the solution of the GreedyML algorithm of a child when it is accessible to the child node.
Next, we state and prove Lemma 4.3 that relates the expected solution of the GreedyML algorithm at a child node with the optimal solution at the node when the approximation ratio of the GreedyML  algorithm at the child is $\beta$.
Lemma 4.3.
Let $c=(\ell-1,id_{c})$ be a child of an internal node $n=(\ell,id)$ of the accumulation tree. Let $S_{c}$ be the solution computed from child $c$, and $V_{c}\subset V_{n}$ denote the elements considered in forming $S_{c}$. If $\mathbb{E}_{V_{c}}[f(S_{c})]\geqslant\beta\cdot f(OPT_{\ell-1,id_{c}})$, then
$\mathop{\mathbb{E}}_{V_{n}}[f(S_{c})]\geqslant\beta\cdot\widehat{f}(1_{OPT_{\ell,id}}-p_{\ell,id}).$
Proof 4.4.
We first construct a subset of $OPT_{\ell,id}$ that contains all the elements that do not appear in $S_{c}$ when added to some leaf node in the subtree rooted at child $c$. Let $O_{c}$ be the rejected set that can be added to $V_{c}$ without changing $S_{c}$; i.e.,
$O_{c}=\{e\in OPT_{\ell,id}:e\notin\textsc{GreedyML}(V_{c}\cup\{e\},\ell^{\prime},id)\}.$
To clarify further, $O_{c}$ is a randomized set dependent on the tape $r_{V_{\ell,id}}$. Since the distribution of $V_{c}$ is the same as $V_{\ell,id}(1/b)$ for each element $e$ in $OPT_{\ell,id}$,
$\Pr[e\in O_{c}]=1-\Pr[e\notin O_{c}]=1-p_{\ell,id}(e).$ (1)
From Lemma 4.1, we know that $\textsc{GreedyML}(V_{c}\cup O_{c},\ell-1,id_{c})=\textsc{GreedyML}(V_{c},\ell-1,id_{c})$ Since the rejected set $O_{c}\subseteq OPT_{\ell,id}$ and the constraints are hereditary, $O_{c}\in\mathcal{C}$ (i.e $O_{c}$ is a feasible solution of child node c).
$f(OPT_{\ell-1,id_{c}})\geqslant f(O_{c}).$ (2)
Then from the condition of Lemma 4.3, we have
$\mathbb{E}_{V_{n}}[\mathbb{E}_{V_{c}}f(S_{c})]$ $\geqslant\beta\cdot\mathbb{E}_{V_{n}}[f(OPT_{\ell-1,id_{c}})]$
$\mathbb{E}_{V_{n}}[f(S_{c})]$ $\geqslant\beta\cdot\mathbb{E}_{V_{n}}[f(OPT_{\ell-1,id_{c}})]$ $[V_{c}\subset V_{n}]$
[…]
Now we show how the solution of the Greedy algorithm that runs at each internal node compares with the optimal solution at the internal node.
Lemma 4.5.
For an internal node $n=(\ell,id)$, let $D$ be the union of all the solutions computed by the children of node $n$ in the accumulation tree. Let $S=\textsc{Greedy}(D)$ be the solution from the Greedy algorithm on the set $D$. If Greedy is an $\alpha$-approximate algorithm, then
$\mathbb{E}_{V_{n}}[f(S)]\geqslant\alpha\cdot f^{-}(p_{\ell,id}).$
Proof 4.6.
We first show a preliminary result on the union set $D$. Consider an element $e\in D\cap OPT_{\ell,id}$ present in some solution $S_{c}$ from a child $c$. Then,
$\Pr[e\in S_{c}|e\in V_{c}]=\Pr[e\in\textsc{GreedyML}(V_{c},\ell-1,id)|e\in V_{c}].$
Since the distribution of $V_{c}\sim V_{\ell,id}(1/b)$ conditioned on $e\in V_{c}$ is identical to the distribution of $B\cup\{e\}$, where $B\sim V_{\ell,id}(1/b)$, we have,
$\Pr[e\in S_{c}|e\in V_{c}]=\Pr_{B\sim V_{\ell,id}(1/b)}[e\in\textsc{GreedyML}(B\cup\{e\},\ell-1,id)]=p_{\ell,id}(e).$
$\Pr(D\cap OPT_{\ell,id})=p_{\ell,id}.$ (3)
$f(S)$ $\geqslant\alpha\cdot f(D\cap OPT_{\ell,id})$
$\mathbb{E}_{V_{n}}[f(S)]$ $\geqslant\mathbb{E}_{V_{n}}[\alpha\cdot f(D\cap OPT_{\ell,id})]$
$\geqslant\alpha\cdot\mathbb{E}_{V_{n}}[f(D\cap OPT_{\ell,id})]$
[…]
Theorem 4.7.
Let $T(m,L,b)$ be an accumulation tree. For a universal set $W$ and random tape $r_{W}$ that maps elements of $W$ to the leaves of the tree $T$, let $V_{\ell,id}\subseteq W$ denote the subset of $W$ accessible to a node $(\ell,id)$. Let $OPT_{\ell,id}$ be an optimal solution computed from the subset $V_{\ell,id}$ for the submodular function $f$ with constraints $\mathcal{C}$. If Greedy is an $\alpha$-approximate algorithm, then
${\mathbb{E}_{V_{\ell,id}}\left[f(\textsc{GreedyML}(V_{\ell,id},\ell,id))\right]\geqslant\dfrac{\alpha}{(\ell+1)}f(OPT_{\ell,id}).}$
Proof 4.8.
We prove this theorem by induction on the level $\ell$.
Base case, $\ell=0$: Here, there is no accumulation step, and we obtain the solution from a single node. Thus we run the Greedy algorithm on $V_{\ell,id}$. The result follows since the Greedy algorithm has the approximation ratio $\alpha$.
Inductive case, $\ell=\ell^{\prime}+1$: We first obtain a relation for the quality of the solutions at level $\ell^{\prime}$ compared to the quality of an optimal solution.
For each child c, let $S_{c}$ be a solution computed by the GreedyML algorithm from the data $V_{c}\subset V_{\ell,id}$;
From the induction hypothesis applied to child $c=(\ell^{\prime},id)$, the approximation ratio obtained as a result of the computation $\textsc{GreedyML}(V_{c},\ell^{\prime},id)$ is $\alpha/(\ell^{\prime}+1)=\alpha/\ell$. This implies that $\mathbb{E}_{V_{l^{\prime},id}}[f(S_{c})]\geqslant\frac{\alpha}{\ell}\cdot f^{-}(1_{OPT_{\ell^{\prime},id}}).$ Therefore we can apply Lemma 4.3 to get
$\mathbb{E}_{V_{\ell,id}}[f(S_{c})]\geqslant\frac{\alpha}{\ell}\cdot\widehat{f}(1_{OPT_{\ell,id}}-p_{\ell,id}).$ (5)
After obtaining the solutions from the children, we get the solution $S$ computed by the Greedy algorithm on the union of these solution sets. From Lemma 4.5, we have
$\mathbb{E}_{V_{\ell,id}}[f(S)]\geqslant\alpha\cdot\widehat{f}(p_{\ell,id}).$ (6)
Now we obtain the relation between the solution at level $\ell^{\prime}+1$ and the optimal solution. Let the solution set at level $\ell^{\prime}+1$ be $T$. We have $T=\arg\max\{f(S),f(S_{c})\}$. Then, we can use the lower bounds calculated earlier in Eqn. 5 and Eqn. 6 to find lower bounds for $T$.
$\mathbb{E}_{V_{\ell,id}}[f(T)]\geqslant\alpha\cdot\widehat{f}(p_{\ell,id})\text{ and }\mathbb{E}_{V_{\ell,id}}[f(T)]\geqslant\frac{\alpha}{\ell}\cdot\widehat{f}(1_{OPT_{\ell,id}}-p_{\ell,id}).$
By multiplying the second inequality by $\ell$ and then adding it to the first, we get
$(\ell+1)\mathbb{E}_{V_{\ell,id}}[f(T)]$ $\geqslant\alpha\cdot(\widehat{f}(1_{OPT_{\ell,id}}-p_{\ell,id})+\widehat{f}(p_{\ell,id}))$
[…]
Dividing by by $\ell+1$, and substituting from Lovász Ext. (1), 2.2 we conclude that the algorithm is $\alpha/(\ell+1)$-approximate.
4.2 Submodular Functions and Complexity
Here, we describe three submodular functions that we consider in our experiments and then discuss their computational and communication complexities.
@@ 4.2 Submodular Functions and Complexity
The first problem we consider is the $k$-cover. Given a ground set B, a collection of subsets $V\subseteq 2^{B}$, and an integer $k$, the goal is to select a set $S\subseteq V$ containing $k$ of these subsets to maximize $f(S)=|\bigcup_{S_{i}\in S}S_{i}|$.
@@
The $k$-dominating set problem is a special case of the $k$-cover problem defined on graphs with the ground set $V$ as the set of vertices. We say a vertex $u\in V$ dominates all its adjacent vertices (denoted by $\delta(u)$). Our goal is to select a set $S$ of $k$ vertices to dominate as many vertices as possible, i.e., $f(S)=|\bigcup_{u\in S}\delta(u)|$. The marginal gain of any vertex is the number of vertices in its neighborhood that are not yet dominated. Therefore the problem shows diminishing marginal gains and is submodular.
@@
The $k$-medoid problem [12] is used to compute exemplar-based clustering, where we want to select a set of exemplars (cluster centers) that are representatives of a large dataset. Given a collection of elements in a ground set $V$, and a dissimilarity $d(u,v)$, we define a loss function (denoted by $L$) as the average pairwise dissimilarity between the exemplars ($S$) and the elements of the data set, i.e., $L(S)=\dfrac{1}{|V|}\sum_{u\in V}\min_{v\in S}d(u,v)$. Following [24], we turn this loss minimization to a submodular maximization problem by setting $f(S)=L(\{e_{0}\}-L(S\cup\{e_{0}\}$, where $e_{0}$ is an auxiliary element specific to the dataset. The goal is to select a subset $S\subseteq V$ of size $k$ representing the exemplars that maximize $f(S)$.
Next, we will analyze the computational and communication complexity of our GreedyML algorithm using the bulk synchronous parallel (BSP) model of parallel computation [29]. For the analysis, we will denote the number of elements in the ground set by $n=|V|,$the solution size by $k$, the number of machines by $m$, and the number of levels in the accumulation tree by $L$.
Computational Complexity. The number of objective function calls by the sequential Greedy algorithm (shown in Algorithm 1 is $O(nk)$, since $k$ elements are selected to be in the solution, and we may need to compute $O(n)$ marginal gains for each of them. Each machine in RandGreedi algorithm makes $O(k(n/m+mk))$ function calls, where the second term comes from the accumulation step. Each machine of the GreedyML algorithm with branching factor $b$ makes $O(k(n/m+Lbk))$ calls. Recall that $L=\lceil\log_{b}{m}\rceil$.

### v3

@@ B.1 Pseudocode of  RandGreedi
B.2 Pseudocode of  GreedyML
@@
Algorithm 2 describes our multilevel distributed algorithm using two procedures. The first procedure GreedyML is a wrapper function that sets up the environment to run the distributed algorithm. The second function $\mbox{\sc GreedyML}^{\prime}$ is the iterative implementation of the recurrence relation that runs on each machine. The wrapper function partitions the data into $m$ subsets and assigns them to the machines (Line 2). Then each machine runs the $\mbox{\sc GreedyML}^{\prime}$ function on the subset assigned to it (Line 5, Line 7). The wrapper function uses and returns the solution from machine 00 (Line 8) as it is the root of the accumulation tree.
@@
The $\mbox{\sc GreedyML}^{\prime}$ procedure is an iterative implementation of the recurrence relation 2 that runs on every machine. Each machine checks whether it needs to be active at a particular level (Line 5) and decides whether it needs to receive from (Line 11) or send to other machines (Line 6). The function returns the solution from the last level of the machine.
@@
3:    for $i=1\ldots m-1\text{ in parallel}$ $\triangleright$ Run GreedyML’ on all machines except 00
@@
5:         Run $\mbox{\sc GreedyML}^{\prime}(V_{i},\ell,b,i)$ to obtain $S_{i}$ on machine $i$
@@
7:    Run $\mbox{\sc GreedyML}^{\prime}(V_{0},\lceil\log_{b}m\rceil,b,0)$ to obtain $S_{0}$ on machine 0
@@
1:procedure $\mbox{\sc GreedyML}^{\prime}$($P$: Partial Data-set, $\ell$: levels; $b$: branching factor, $id$: machine ID)
@@
Appendix C Submodular Functions and Complexity
@@
Given a ground set $B$, a collection of subsets $V\subseteq 2^{B}$, and an integer $k$, the goal is to select a set $S\subseteq V$ containing $k$ of these subsets to maximize $f(S)=|\bigcup_{S_{i}\in S}S_{i}|$.
@@
The $k$-dominating set problem is a special case of the $k$-cover problem defined on graphs with the ground set $V$ as the set of vertices. We say a vertex $u\in V$ dominates all its adjacent vertices (denoted by $\delta(u)$). Our goal is to select a set $S$ of $k$ vertices to dominate as many vertices as possible, i.e., $f(S)=|\bigcup_{u\in S}\delta(u)|$. The marginal gain of any vertex is the number of vertices in its neighborhood that are not yet dominated. Therefore, the problem shows diminishing marginal gains and is submodular.
@@
The $k$-medoid problem [10] is used to compute exemplar-based clustering, which asks for a set of exemplars (cluster centers) representatives of a large dataset. Given a collection of elements in a ground set $V$, and a dissimilarity measure $d(u,v)$, we define a loss function (denoted by $L$) as the average pairwise dissimilarity between the exemplars ($S$) and the elements of the data set, i.e., $L(S)=\nicefrac{{1}}{{|V|}}\sum_{u\in V}\min_{v\in S}d(u,v)$. Following [21], we turn this loss minimization to a submodular maximization problem by setting $f(S)=L(\{e_{0}\}-L(S\cup\{e_{0}\}$, where $e_{0}$ is an auxiliary element specific to the dataset. The goal is to select a subset $S\subseteq V$ of size $k$ that maximizes $f(S)$.
Next, we analyze the computational and communication complexity of our GreedyML algorithm using the bulk synchronous parallel (BSP) model of parallel computation [28]. We denote the number of elements in the ground set by $n=|V|,$the solution size by $k$, the number of machines by $m$, and the number of levels in the accumulation tree by $L$.
Computational Complexity. The number of objective function calls by the sequential Greedy algorithm is $O(nk)$, since $k$ elements are selected to be in the solution, and we may need to compute $O(n)$ marginal gains for each of them. Each machine in RandGreedi algorithm makes $O(k(n/m+mk))$ function calls, where the second term comes from the accumulation step. Each machine of the GreedyML algorithm with branching factor $b$ makes $O(k(n/m+Lbk))$ calls. Recall that $L=\lceil\log_{b}{m}\rceil$.